### v2

# HERO: Heterogeneous Continual Graph Learning via Meta-Knowledge Distillation

Abstract.
Heterogeneous graph neural networks have seen rapid progress in web applications such as social networks, knowledge graphs, and recommendation systems, driven by the inherent heterogeneity of web data. However, existing methods typically assume static graphs, while real-world graphs are continuously evolving. This dynamic nature requires models to adapt to new data while preserving existing knowledge. To this end, this work introduces HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation), a unified framework for continual learning on heterogeneous graphs. HERO employs meta-adaptation, a gradient-based meta-learning strategy that provides directional guidance for rapid adaptation to new tasks with limited samples. To enable efficient and effective knowledge reuse, we propose DiSCo (Diversity Sampling with semantic Consistency), a heterogeneity-aware sampling method that maximizes target node diversity and expands subgraphs along metapaths, retaining critical semantic and structural information with minimal overhead. Furthermore, HERO incorporates heterogeneity-aware knowledge distillation, which aligns knowledge at both the node and semantic levels to balance adaptation and retention across tasks. Extensive experiments on four web-related heterogeneous graph benchmarks demonstrate that HERO substantially mitigates catastrophic forgetting while achieving efficient and consistent knowledge reuse in dynamic web environments.
Incremental Learning, Model Reuse, Stability-Plasticity Trade-off
Corresponding authors: Xikun Zhang and Dongjin Song
$c$s: Computing methodologies Lifelong machine learningccs: Information systems Social networks
1. Introduction
The modern web is inherently heterogeneous, consisting of diverse entities and relations such as users, items, hyperlinks, and knowledge facts. As a result, heterogeneous graphs (HGs) have become a natural abstraction for modeling complex web systems including social networks, knowledge graphs, and recommendation platforms (Shi et al., 2018; Sun and Han, 2012). Unlike homogeneous graphs with uniform node and edge types, HGs encode multifaceted relationships across different types of nodes and edges, enabling rich and expressive representations for web-scale applications. For example, in recommendation systems, user–item interactions form user nodes, item nodes, and purchase edges; in knowledge graphs, entities and relations compose semantic networks that support web search and reasoning; in social media, dynamic interactions between users, posts, and topics constitute evolving heterogeneous graphs that drive personalized content delivery. To extract meaningful insights from heterogeneous graphs, different Heterogeneous Graph Neural Networks (HGNNs) have been developed. These approaches mainly fall into two categories: metapath-based HGNNs (Fu et al., 2020; Zhang et al., 2019; Wang et al., 2019; Schlichtkrull et al., 2018; Dong et al., 2017; Sun et al., 2011; Yun et al., 2019; Yang et al., 2023; Zhu et al., 2024), which aggregate information over predefined meta-paths to mine the intricate relationship among the heterogeneous nodes and edges, and meta-path-free models (Hu et al., 2020; Lv et al., 2021; Zhu et al., 2019; Hong et al., 2020; Zhang et al., 2022e; Fu et al., 2023; Huo et al., 2025) that automatically learn node interactions without preconfigured paths, offering enhanced adaptability but potentially sacrificing interpretability.
Figure 1. An example of Evolving Heterogeneous Graph. (a) Node types include Actor, Movie, and Director. (b) As the graph evolves, new domains emerge (e.g., movies of different genres such as Comedy/Romance and Thriller/Action).
Despite the success of Heterogeneous Graph Neural Networks (HGNNs) in static settings, real-world web graphs are inherently evolving. In movie-related graphs such as IMDB, new entities and relations emerge over time, reflecting shifts in genres, collaborations, and audience interests (Figure 1). In such evolving environments, HGNNs are expected to continuously incorporate the incoming new knowledge while preserving the learnt patterns. However, existing continual learning (CL) methods, largely developed for images (Aljundi et al., 2018; Kirkpatrick et al., 2017; Li and Hoiem, 2017; Lopez-Paz and Ranzato, 2017) and homogeneous graphs (Zhang et al., 2022a; Liu et al., 2021; Zhou and Cao, 2021; Zhang et al., 2023b; Unal et al., 2023; Su et al., 2023; Ren et al., 2023; Niu et al., 2024; Li et al., 2024; Mondal et al., 2024; Liu et al., 2024; Choi et al., 2024; Qiao et al., 2025), fall short when directly applied to heterogeneous graphs due to semantic diversity and structural imbalance. For instance, in recommendation platforms, rare product categories are easily forgotten; in knowledge graphs, semantically distinct relations collapse into uniform representations; and in social networks, global regularization fails to capture type-specific parameter importance. These challenges suggest the necessity to develop new continual learning strategies which can explicitly account for web-scale heterogeneity.
To bridge this critical gap, we present a systematic investigation of the Heterogeneous Continual Graph Learning (HCGL) problem and introduce HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation), a unified framework for continual learning on heterogeneous graphs. Unlike homogeneous graphs, the diverse node and edge types in heterogeneous graph introduce extra challenges to knowledge transfer across different tasks. Therefore, for effective model adaptation over the continuously evolving heterogeneous graph structures, we design a Gradient-based Meta-learning Module (G-MM). G-MM learns an optimized initialization of model parameters by leveraging gradient-based meta-learning, which allows fast model adaptation to new tasks using only a small number of examples. Through task-specific adjustments, G-MM also avoids the task-wise interference and ensures performance robustness on learned tasks. Memory replay has proven to be an effective strategy for mitigating catastrophic forgetting. However, existing replay methods typically rely on storing large volumes of historical data, which makes continual learning impractical in large-scale tasks. Moreover, conventional memory sampling strategies cannot be directly applied to HCGL due to the complexity of heterogeneous structures. To address this, we propose DiSCo (Diversity Sampling with semantic Consistency), a heterogeneity-aware sampling method that maximizes the diversity of target nodes and expands subgraphs along meta-paths, thereby retaining critical semantic and structural information with minimal memory overhead. Furthermore, to align knowledge across tasks and maintain semantic consistency in heterogeneous settings, we introduce a Heterogeneity-aware Knowledge Distillation (HKD) module. Beyond knowledge retained via experience replay, this module performs both logit-level and semantic-level alignment between the current and previous task models. By distilling prediction distributions and meta-path-based attention representations, HKD guides the student model to retain predictive information while incorporating multi-level semantic information, significantly improving continual learning performance on heterogeneous graphs. Our contributions are summarized as follows:
  1. (1)
    We provide the first systematic investigation of the Heterogeneous Continual Graph Learning (HCGL) problem, highlighting its unique challenges in web-scale settings compared to homogeneous graphs.
  2. (2)
    We propose DiSCo, a diversity- and semantics-aware sampling strategy that enables memory-efficient and structurally balanced experience replay.
  3. (3)
    We design a heterogeneity-aware knowledge distillation mechanism that aligns knowledge at both logit and semantic levels to balance adaptation and retention.
  4. (4)
    We develop the HERO framework, a novel framework tailored for HCGL. Experiments on multiple heterogeneous web graph benchmarks (e.g., knowledge graph, recommendation and citation networks datasets) validate its superior performance in terms of accuracy, efficiency, and robustness.
2. Related Work
Heterogeneous Graph Neural Networks.
Heterogeneous graphs are prevalent in web environments such as knowledge graphs, social networks, and recommendation systems, where multiple node and edge types naturally coexist. These graphs are characterized by complex topological structures and rich semantic information, which traditional Graph Neural Networks (GNNs) struggle to model directly. To address this challenge, specialized approaches known as Heterogeneous Graph Neural Networks (HGNNs) have been developed to effectively capture the multi-typed relationships inherent in web graphs. Existing HGNN methods can be broadly categorized into two groups: metapath-based and metapath-free approaches. Metapath-based methods (Fu et al., 2020; Zhang et al., 2019; Wang et al., 2019; Schlichtkrull et al., 2018; Dong et al., 2017; Sun et al., 2011; Yun et al., 2019; Yang et al., 2023; Zhu et al., 2024) explicitly leverage predefined semantic paths, aggregating features from neighbors within the same semantic context and integrating across metapaths. In contrast, metapath-free methods (Hu et al., 2020; Lv et al., 2021; Zhu et al., 2019; Hong et al., 2020; Zhang et al., 2022e; Fu et al., 2023; Huo et al., 2025) adopt a more flexible design by aggregating information from all types of neighbors within a local neighborhood, while encoding semantic distinctions (e.g., node and edge types) through mechanisms such as attention. These methods have achieved remarkable success on static web graphs; however, they typically assume full access to the entire graph during training. This assumption becomes restrictive in realistic web scenarios, where graphs evolve continuously with new nodes and relations. To this end, Continual Learning (CL) for HGNNs emerges as a crucial yet underexplored research direction, enabling models to adapt to dynamic web graphs while preserving previously acquired knowledge.
Continual Graph Learning.
Continual Graph Learning (CGL) enables models to learn from evolving graph-structured data while mitigating catastrophic forgetting, where performance on earlier tasks deteriorates after learning new ones. Existing methods can be broadly categorized into three types: parameter-isolation, regularization, and memory-replay approaches. Parameter-isolation methods preserve important parameters by freezing them after training on a task, preventing interference from subsequent updates. Examples include HPNs (Zhang et al., 2022b) and the PI-GNN framework (Zhang et al., 2023b). Regularization methods constrain updates to important parameters via penalty terms. Representative methods include EWC (Kirkpatrick et al., 2017), MAS (Aljundi et al., 2018), and TWP (Liu et al., 2021), which considers local graph topology. Knowledge distillation is also commonly used to retain information from prior models (Xu et al., 2020; Dong et al., 2021). Memory-replay methods sample and store representative nodes or subgraphs for replay (Zhang et al., 2023a; Ahrabian et al., 2021). However, these methods may suffer from memory explosion due to the growth of computational subgraphs. ER-GNN (Zhou and Cao, 2021) addresses this by sampling node attributes only, while SSM (Zhang et al., 2022c) stores sparsified subgraphs. DSLR (Choi et al., 2024) further enhances replay by promoting sample diversity and learning informative subgraph structures to better preserve past knowledge. PDGNNs-TEM (Zhang et al., 2024b) maintains dynamic embeddings of subgraphs, and UGCL (Hoang et al., 2023) unifies memory replay and distillation to support both node and graph classification tasks. TACO (Han et al., 2024) replaces raw subgraph storage with graph coarsening, enabling compact memory usage while preserving key structural information. A recent work, RL-GNN (Song et al., 2025), addresses graph-level continual learning by identifying invariant rationales across tasks to mitigate catastrophic forgetting caused by spurious correlations. Despite recent progress, CGL still faces challenges in reducing forgetting and computational cost. Moreover, continual learning on heterogeneous graphs remains underexplored. In this work, we address these challenges by using proposed DiSCo and Heterogeneity-aware Knowledge Distillation tailored to heterogeneous graph settings.
Meta Learning.
Meta-learning aims to enable models to adapt rapidly to new tasks by leveraging shared knowledge. Existing approaches are generally categorized into optimization-based (e.g., MAML (Finn et al., 2017), Reptile (Nichol and Schulman, 2018)) and metric-based methods (e.g., Prototypical Networks (Snell et al., 2017), Matching Networks (Vinyals et al., 2016)). By learning how to learn, meta-learning provides a principled way to capture transferable inductive biases that facilitate fast adaptation to new tasks with limited data. In the context of continual learning, meta-learning has been applied to both online (Gupta et al., 2020) and few-shot settings (Javed and White, 2019), where it helps alleviate catastrophic forgetting by enabling rapid re-optimization when task distributions shift. Recently, the integration of meta-learning with graph neural networks has drawn increasing attention, as graphs naturally exhibit diverse and evolving structures that demand adaptive learning capabilities. For example, MetaCLGraph (Unal et al., 2023) incorporates meta-learning with experience replay to balance adaptation and stability across sequential graph tasks, while HG-Meta (Zhang et al., 2022d) extends meta-learning to few-shot scenarios on heterogeneous graphs by modeling task-level semantics and relational dependencies. Motivated by these advances, we further explore the potential of meta-learning in mitigating forgetting and improving adaptability under heterogeneous continual graph learning, where both structural and semantic variations pose additional challenges compared to conventional settings.
3. Preliminary
In this section, we introduce fundamental concepts and formalize the problem of Heterogeneous Continual Graph Learning (HCGL). Our primary objective is to develop a continual learning framework on Evolving Heterogeneous Graphs to mitigate the issue of catastrophic forgetting. In the HCGL setting, a Heterogeneous Graph Neural Network (HGNN) learns a sequence of tasks in a domain incremental setting (Zhang et al., 2024a) (See Appendix B.3 for the details of this setting), without access to the data from previous tasks. However, it is allowed to utilize a memory buffer with limited capacity to store representative information. The goal of the framework is to maximize the prediction accuracy across all tasks after training while minimizing the forgetting of previously acquired knowledge during the learning of new tasks.
In this work, we focus on node classification tasks and adopt a common continual learning setup in which the dataset is partitioned into a sequence of tasks based on node class labels. In this way, different splits have non-overlapping category labels.
Definition 1 (Heterogeneous Graph).
A heterogeneous graph is defined as $G=(V,E)$, where $V$ is the set of nodes, and $E$ is the set of edges. Each node $v\in V$ is associated with a type given by the node-type mapping function $\phi:V\rightarrow\mathcal{A}$, and each edge $e\in E$ is associated with a type given by the edge-type mapping function $\psi:E\rightarrow\mathcal{R}$, where $\mathcal{A}$ and $\mathcal{R}$ denote the sets of node and edge types, respectively. The graph is considered heterogeneous if $|\mathcal{A}|+|\mathcal{R}|>2$. Let $V_{\tau}$ denote the set of nodes of type $\tau\in\mathcal{T}$.
Definition 2 (Metapath).
A metapath $\mathcal{P}$ is a sequence of node and edge types in the form of $A_{1}\xrightarrow{R_{1}}A_{2}\xrightarrow{R_{2}}\cdots\xrightarrow{R_{l}}A_{l+1}$, which defines a composite relation $R_{1}\circ R_{2}\circ\cdots\circ R_{l}$ between node types $A_{1}$ and $A_{l+1}$, where $\circ$ denotes the composition operator. The metapath captures high-level semantic connections across different types of nodes through a specific sequence of edge types.
Definition 3 (Heterogeneous Continual Graph Learning).
Let $\mathcal{G}=(V,E)$ be a dynamic heterogeneous graph, where $V$ is the node set and $E$ is the edge set. The feature set is type-specific, i.e., $\mathcal{F}=\{F_{\tau}\in\mathbb{R}^{|V_{\tau}|\times d_{\tau}}\mid\tau\in\mathcal{A}\}$. In the HCGL setting, the model learns a sequence of tasks $\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$, without access to the data from previous tasks. Specifically, each task $\mathcal{T}_{t}$ corresponds to a subgraph $\mathcal{G}_{t}$ with disjoint category labels, and due to storage limitations, the model can only access the data available at the current task. For each subgraph $\mathcal{G}_{t}$, we divide it into a training set $\mathcal{G}_{t}^{\text{tr}}=(V_{t}^{\text{tr}},E_{t}^{\text{tr}})$ and a testing set $\mathcal{G}t^{\text{te}}=(V_{t}^{\text{te}},E_{t}^{\text{te}})$ to train and evaluate the model $f_{\theta}$.
4. Methodology
In this section, we present our HERO (HEterogeneous continual gRaph learning via meta-knOwledge distillation) framework (see Figure 2), provide a detailed introduction of its core components, and explain how it can systematically address the identified challenges.
Figure 2. The overall framework of HERO. (a) Diversity Sampling with semantic Consistency: Construct task-specific subgraphs by selecting diverse target-type nodes and expanding to related node types via relation-aware importance. (b) Heterogeneity-aware Knowledge Distillation: Align previous and current tasks via logit-level and semantic-level distillation. (c) Task Training: Use meta-learning for fast adaptation, jointly optimizing task loss, replay loss, and distillation loss.
4.1. Fast Adaptation with Gradient-based Meta-Learning
Efficient adaptation to new data patterns is crucial for ensuring the practical usability of a model in an evolving heterogeneous web graph. To achieve this, we introduce the Gradient-based Meta-learning Module (G-MM), which optimizes model parameter initialization to enable rapid adaptation to new tasks. This, in turn, helps preserve performance on the current task. Specifically, given the training node set $V_{t}^{tr}$ of the current task $\mathcal{T}_{t}$, we select $e$ samples from $V_{t}^{tr}$ based on the Coverage Maximization (CM) strategy (introduced in Section 4.2), which identifies key nodes by maximizing node diversity. In practice, labeled data in heterogeneous graphs is often highly limited. By performing gradient descent on the selected small sample set $\mathcal{E}$, the model can quickly adapt to the current task, a process referred to as the inner update. Let $\theta$ denote the set of model parameters. For the current task $\mathcal{T}_{i}$, We feed the sampled node set $\mathcal{E}$ into the model and compute the loss $\mathcal{L}_{\mathcal{E}}$, updating $\theta$ to $\theta_{fast}$ via gradient descent (Inner Update):
(1) $\theta_{fast}\leftarrow\theta-\alpha\nabla_{\theta}\mathcal{L}_{\mathcal{E}}(\theta),$
where $\alpha$ is the step size for the inner update. Meta-learning with a small number of samples enables the model to prevent overfitting during experience replay while ensuring rapid adaptation to the current task, thereby maintaining robust overall performance.
4.2. Diversity Sampling with semantic Consistency
In addition to adapting to new data patterns, overcoming catastrophic forgetting is crucial, as previously observed patterns may reappear in practical scenarios. While many existing experience replay methods primarily consider homogeneous graphs or single-type nodes, they fail to capture the complex semantic and structural dependencies in heterogeneous graphs. To address this, we propose DiSCo (Diversity Sampling with semantic Consistency), which jointly considers the diversity of nodes within the target node type and the structural connections with nodes of other types. DisCo proceeds in two stages: (1) selecting representative target-type nodes by maximizing diversity in the feature or embedding space, and (2) expanding to other types of nodes through relation-type-aware importance estimation, ensuring that both semantic diversity and heterogeneous structural patterns are preserved.
Step 1: Target Node Selection via Coverage Maximization. Prior research has shown that different data points contribute unequally to model training: some significantly enhance performance, while others offer limited benefit (Toneva et al., 2018). Motivated by this, we identify key nodes from the current task by maximizing node diversity using the Coverage Maximization (CM) strategy. This approach allows us to retain essential historical information with only a small number of nodes. Formally, given the training node set $V_{t}^{tr}$ of task $\mathcal{T}_{t}$, CM selects a subset by maximizing the coverage of the attribute/embedding space:
(2) $\mathcal{N}(v_{i})=\{v_{j}\mid\text{dist}(v_{i},v_{j})<d,\,\mathcal{Y}(v_{i})\neq\mathcal{Y}(v_{j})\},$
where $\mathcal{Y}(v_{i})$ denotes the label of node $v_{i}$, and $\mathcal{N}(v_{i})$ represents the set of nodes from different classes that are within a distance $d$ of $v_{i}$. Given the memory buffer $\mathcal{B}_{\tau_{t}}$ for the target node type $\tau_{t}$, we select $e$ nodes per class with the smallest $|\mathcal{N}(v_{i})|$ as representative experiences.
Step 2: Multi-hop Neighbor Expansion. Given target nodes $V_{\tau_{t}}$, we collect candidate nodes $\mathcal{C}_{r}$ for each relation type $r\in\mathcal{R}$ through:
(3) $\mathcal{C}_{r}=\bigcup_{v\in V_{\tau_{t}}}\left\{u\,\middle|\,(v,u)\in E_{r}\lor(u,v)\in E_{r}\right\},$
where $E_{r}$ denotes edges of relation type $r$. We define node importance $\pi(v)$ as the sum of relation-specific degrees:
(4) $\pi(v)=\sum_{r\in\mathcal{R}}\deg_{r}(v),$
where $\deg_{r}(v)$ measures the connectivity strength of node $v$ under relation type $r$. For each node type $\tau$ with buffer $\mathcal{B}_{\tau}$, we select top-$|\mathcal{B}_{\tau}|$ nodes by $\pi(v)$:
(5) $V^{*}_{\tau}=\text{topk}_{v\in\mathcal{C}_{\tau}}\left(\pi(v),|\mathcal{B}_{\tau}|\right),$
With sampled target type node set $V_{\tau_{t}}$ and the selected heterogeneous neighbor nodes set $\{V_{\tau}\}_{\tau\in\mathcal{A}/\tau_{t}}$, we construct the heterogeneous subgraph $\mathcal{G}_{sub}=(V^{\prime},E^{\prime})$ through:
(6) $V^{\prime}=V_{\tau_{t}}\cup\bigcup_{\tau\neq\tau_{t}}V_{\tau},\quad E^{\prime}=\{(u,v)\in E\mid u,v\in V^{\prime}\},$
For all $v\in V_{\tau_{t}}$, we retain their original features and labels, and maintain the structure of the type-specific projection. When learning a new task, we perform the experience replay by incorporating an auxiliary loss computed over the memory buffer $\mathcal{B}$ into the current task loss (i.e., the cross-entropy loss between the given labels $\mathcal{Y}$ and the predicted labels $\hat{\mathcal{Y}}$)
(7) $\mathcal{L}$ $=\mathcal{L}_{\mathcal{T}_{t}}(\mathcal{G}_{t},\theta)+\lambda_{er}\mathcal{L}_{er}(\mathcal{G}_{sub},\theta)$
$=-\sum_{v_{i}\in\mathcal{V}_{t}^{tr}}\mathcal{Y}(v_{i})\log\hat{\mathcal{Y}}(v_{i})-\lambda_{er}\sum_{v_{j}\in\mathcal{B}}\mathcal{Y}(v_{j})\log\hat{\mathcal{Y}}(v_{j}),$
where $\hat{\mathcal{Y}}(v_{i})$ is the predicted label of node $v_{i}$, $\lambda_{er}$ modulates the contribution of buffered data in the overall loss.
4.3. Heterogeneity-aware Knowledge Distillation for Task Alignment
Existing knowledge distillation methods are often developed for homogeneous scenarios and fail to account for the multi-relational semantics and cross-type dependencies in heterogeneous graphs. In contrast, our Heterogeneity-aware Knowledge Distillation (HKD) module is explicitly tailored for heterogeneous graph continual learning. Instead of relying solely on logit-based distillation, we propose a two-level alignment strategy that leverages both prediction and semantic signals to capture graph heterogeneity. The logit-level distillation transfers soft label distributions to retain discriminative knowledge, while the semantic-level alignment matches meta-path-aware attention scores, enabling the student model to preserve high-order structural patterns unique to heterogeneous graphs.
Logit-level Distillation Inspired by previous distillation-based approaches (Tian et al., 2023), we transfer the soft knowledge from the teacher model (previous task classifier) to the student model (current task classifier), enabling the student model to learn both new knowledge from the current task and retained knowledge from past tasks. Mimicking teacher’s prediction results enables the student model to learn the secondary information from previous tasks that cannot be expressed by the experience replay data alone. Soft knowledge from teacher model is formulated as the predicted probability of the labels in the current task data:
(8) $P^{T}(z_{i},t)=\text{Softmax}\big(f_{T}(z_{i}),t\big)=\frac{\exp\left[f_{T}(z_{i})/t\right]}{\sum_{j}\exp\left[f_{T}(z_{j})/t\right]},$
where $z_{i}$ is the embedding of node $v_{i}$ in $\mathcal{V}_{i}^{tr}$, $f_{T}(z_{i})$ is the score logit of $z_{i}$ obtained from teacher model, and $t$ is the temperature index to soften the peaky softmax distribution (Hinton, 2015). Thus, the knowledge distillation loss of the teacher model and the student model is defined as follows:
(9) $\mathcal{L}_{\text{logit}}=\text{Mean}\big(t^{2}\sum_{l}^{N}\sum_{v_{i}\in\mathcal{V}_{i}^{tr}}P_{l}^{T}(z_{i},t)\log\frac{P_{l}^{T}(z_{i},t)}{P^{S}(z_{i},t)}\big),$
where $N$ is the number of teacher models, $P^{T}$ and $P^{S}$ are the predicted distributions of teacher model and student model respectively.
Semantic-level Distillation To preserve metapath-induced semantic patterns, we propose an attention-based structural distillation method. Given a predefined metapath set $\mathcal{P}=\{P_{1},\dots,P_{M}\}$, let $\boldsymbol{\alpha}_{P_{m}}^{(T)}(v_{i})$ and $\boldsymbol{\alpha}_{P_{m}}^{(S)}(v_{i})$ denote the attention coefficients for metapath $P_{m}$ computed by the teacher and student models, respectively. These attention coefficients reflect the importance of different semantic contexts encoded by heterogeneous structural patterns. To align the semantic-level knowledge between the teacher and the student, we define the semantic alignment loss as:
(10) $\mathcal{L}_{\text{sem}}=\sum_{m=1}^{M}\|\boldsymbol{\alpha}_{P_{m}}^{(T)}-\boldsymbol{\alpha}_{P_{m}}^{(S)}\|_{2}$
where $\boldsymbol{\alpha}_{P_{m}}\in\mathbb{R}^{|V^{tr}|}$ is the normalized attention vector over all nodes for metapath $P_{m}$}. We provide a detailed discussion in Appendix A on how this module can be extended to arbitrary HGNNs beyond metapath-based architectures.
Algorithm 1 Details of the HERO Framework.
1:Input: Continual tasks $\mathcal{T}=\{\mathcal{T}_{1},\mathcal{T}_{2},\dots,\mathcal{T}_{T}\}$; Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$; Learning rate $\alpha$ for the inner update; Learning rate $\beta$; Number of target type nodes in each class added to $\mathcal{B}$: e.
2:Output: Model $f_{\theta}$ which can mitigate catastrophic forgetting of preceding tasks.
3:Initialize model parameters $\theta$
4:while Continual Tasks $\mathcal{T}$ remains do
5:  Obtain current training set $V^{tr}_{t}$
6:  Obtain teacher models $\{f_{T}\}$ and Buffer set $\{\mathcal{B}_{\tau}\}_{\tau\in\mathcal{A}}$
7:  Select meta-learning examples $\mathcal{E}=Select(\mathcal{V}_{i}^{tr},e)$
8:  for $epoch=1$ to $E$ do
9:   Update the parameters $\theta$ using $\mathcal{E}$ via Equation 1
10:   Replay the experience via Equation 7
11:   Distill the soft knowledge and semantic information from teacher models via Equation 9 and 10
12:   Update the parameters of student model by optimizing Equation 12
13:  end for
14:  Select target type nodes via 2 and other types nodes via 5
15:  Add all sampled nodes to Buffer
16:  Add current task model to teacher models $\{f_{T}\}$
17:  $\mathcal{T}=\mathcal{T}\setminus\{\mathcal{T}_{i}\}$
18:end while
19:Return Model $f_{\theta}$
This loss guides the student model to capture relational importance and structural dependencies aligned with the teacher model, thereby preserving semantic information that is critical in heterogeneous graphs. Afterwards, we combine the logit level loss and the semantic level loss:
(11) $\mathcal{L}_{kd}=\lambda_{\text{logit}}\mathcal{L}_{\text{logit}}+\lambda_{\text{sem}}\mathcal{L}_{\text{sem}},$
where $\lambda_{\text{logit}}$ and $\lambda_{\text{sem}}$ are both trade-off weights for balancing the losses. The final objective of node representation learning is to minimize the joint loss including current task loss $\mathcal{L}_{\mathcal{T}_{i}}$, experience replay loss $\mathcal{L}_{er}$, and the KD loss $\mathcal{L}_{kd}$:
(12) $\mathcal{L}_{joint}=\mathcal{L}_{\mathcal{T}_{t}}+\lambda_{er}\mathcal{L}_{er}+\lambda_{kd}\mathcal{L}_{kd},$
where $\lambda_{kd}$ is a trade-off weight for balancing the losses. Taking the parameters $\theta_{fast}$ obtained from Equation 1 as initialization, the student model for the current task is further updated as follows:
(13) $\theta^{*}=\theta-\beta\nabla_{\theta}\mathcal{L}_{joint}(\theta),$
where $\beta$ is learning rate. By minimizing $\mathcal{L}_{joint}$, parameters $\theta$ of HGNN model are optimized for downstream node classification task. As the number of tasks grows, the teacher models would increase linearly. For scalability, we adapt a sliding window strategy that only keeps the latest three teacher models for knowledge distillation. The details of the HERO framework is provided in 1.
Table 1. Performance comparison with different HGNN backbones on three benchmark datasets. The symbol $\uparrow$ ($\downarrow$) means higher (lower) is better. The best results are highlighted in bold, while the second best results are underlined. ”OOM” means that the model runs out of memory on large graphs.
Base Models Methods DBLP IMDB Freebase
   AP / % $\uparrow$ AF / % $\downarrow$ AP / % $\uparrow$ AF / % $\downarrow$ AP / % $\uparrow$ AF / % $\downarrow$
HAN Finetune 82.8 $\pm$ 4.6 26.6 $\pm$ 9.1 68.8 $\pm$ 2.2 26.0 $\pm$ 2.8 53.8 $\pm$ 4.7 25.7 $\pm$ 8.8
EWC (Kirkpatrick et al., 2017) 85.4 $\pm$ 5.5 21.5 $\pm$ 10.6 69.6 $\pm$ 1.2 25.8 $\pm$ 2.4 58.3 $\pm$ 2.4 18.4 $\pm$ 4.2
MAS (Aljundi et al., 2018) 91.1 $\pm$ 0.8 8.3 $\pm$ 1.5 72.0 $\pm$ 2.3 20.5 $\pm$ 3.9 59.0 $\pm$ 1.0 14.8 $\pm$ 0.9
TWP (Liu et al., 2021) 90.5 $\pm$ 1.3 10.1 $\pm$ 2.6 74.8 $\pm$ 2.7 14.5 $\pm$ 4.4 60.9 $\pm$ 4.3 10.8 $\pm$ 6.9
ER-GNN (Zhou and Cao, 2021) 89.7 $\pm$ 2.9 13.1 $\pm$ 6.1 74.9 $\pm$ 2.3 12.7 $\pm$ 4.2 56.5 $\pm$ 1.3 19.6 $\pm$ 0.9
MetaCLGraph (Unal et al., 2023) 89.9 $\pm$ 0.9 11.4 $\pm$ 2.3 74.8 $\pm$ 3.2 14.8 $\pm$ 5.7 59.4 $\pm$ 2.5 13.0 $\pm$ 4.1
FTF-ER (Pang et al., 2024) 90.8 $\pm$ 1.4 9.8 $\pm$ 2.4 72.0 $\pm$ 4.2 19.7 $\pm$ 6.4 58.1 $\pm$ 0.7 14.6 $\pm$ 0.6
Ours 93.6 $\pm$ 0.8 4.3 $\pm$ 1.6 78.4 $\pm$ 1.7 7.3 $\pm$ 3.0 60.9 $\pm$ 1.5 11.1 $\pm$ 1.6
Joint Train 95.2 $\pm$ 0.6 1.7 $\pm$ 1.0 80.4 $\pm$ 0.7 3.9 $\pm$ 0.8 65.6 $\pm$ 0.9 2.8 $\pm$ 2.0
HGT Finetune 86.2 $\pm$ 1.5 12.7 $\pm$ 2.2 75.1 $\pm$ 7.5 17.0 $\pm$ 9.4 67.8 $\pm$ 1.8 16.5 $\pm$ 2.5
EWC (Kirkpatrick et al., 2017) 89.1 $\pm$ 1.2 10.7 $\pm$ 1.8 77.3 $\pm$ 0.4 12.5 $\pm$ 0.2 69.6 $\pm$ 3.0 14.7 $\pm$ 2.9
MAS (Aljundi et al., 2018) 89.8 $\pm$ 2.5 9.0 $\pm$ 3.0 77.5 $\pm$ 1.2 11.8 $\pm$ 2.6 70.4 $\pm$ 3.1 14.8 $\pm$ 2.0
TWP (Liu et al., 2021) 89.1 $\pm$ 2.1 10.3 $\pm$ 1.1 77.7 $\pm$ 0.6 11.1 $\pm$ 1.6 69.6 $\pm$ 2.0 15.1 $\pm$ 0.9
ER-GNN (Zhou and Cao, 2021) 90.7 $\pm$ 1.5 3.5 $\pm$ 0.4 77.3 $\pm$ 2.2 9.0 $\pm$ 3.6 69.5 $\pm$ 1.4 13.6 $\pm$ 1.7
MetaCLGraph (Unal et al., 2023) 90.0 $\pm$ 0.7 10.9 $\pm$ 1.4 76.1 $\pm$ 0.3 12.2 $\pm$ 1.8 69.1 $\pm$ 2.2 16.3 $\pm$ 2.5
FTF-ER (Pang et al., 2024) 89.4 $\pm$ 2.2 5.5 $\pm$ 1.9 76.1 $\pm$ 1.4 8.9 $\pm$ 3.5 69.7 $\pm$ 2.8 13.0 $\pm$ 2.5
Ours 92.8 $\pm$ 0.8 3.3 $\pm$ 1.2 78.4 $\pm$ 0.4 6.6 $\pm$ 1.7 70.6 $\pm$ 1.2 9.1 $\pm$ 2.7
Joint Train 93.9 $\pm$ 0.6 1.7 $\pm$ 1.0 80.0 $\pm$ 0.7 2.0 $\pm$ 0.9 72.5 $\pm$ 0.6 8.5 $\pm$ 1.2
MAGNN Finetune 79.1 $\pm$ 5.1 34.8 $\pm$ 10.1 71.3 $\pm$ 0.6 21.7 $\pm$ 3.0 OOM OOM
EWC (Kirkpatrick et al., 2017) 91.0 $\pm$ 2.3 10.9 $\pm$ 4.5 73.4 $\pm$ 1.6 16.7 $\pm$ 2.5 OOM OOM
MAS (Aljundi et al., 2018) 92.1 $\pm$ 2.1 7.8 $\pm$ 5.3 74.3 $\pm$ 2.0 14.2 $\pm$ 3.7 OOM OOM
TWP (Liu et al., 2021) 91.6 $\pm$ 2.1 7.4 $\pm$ 5.0 74.0 $\pm$ 0.8 13.7 $\pm$ 3.3 OOM OOM
ER-GNN (Zhou and Cao, 2021) 90.0 $\pm$ 2.1 12.4 $\pm$ 4.3 75.3 $\pm$ 0.9 10.3 $\pm$ 1.5 OOM OOM
MetaCLGraph (Unal et al., 2023) 92.3 $\pm$ 0.4 2.1 $\pm$ 0.7 75.6 $\pm$ 1.0 11.8 $\pm$ 1.3 OOM OOM
FTF-ER (Pang et al., 2024) 92.8 $\pm$ 0.8 1.9 $\pm$ 2.3 73.0 $\pm$ 0.6 15.8 $\pm$ 1.4 OOM OOM
Ours 93.2 $\pm$ 0.5 3.1 $\pm$ 0.8 76.8 $\pm$ 1.1 7.5 $\pm$ 1.4 OOM OOM
Joint Train 94.1 $\pm$ 0.1 1.9 $\pm$ 1.0 77.4 $\pm$ 0.7 4.7 $\pm$ 1.5 OOM OOM
5. Experiments
We conducted experiments on four widely used web graph benchmarks: DBLP (Lv et al., 2021), IMDB (Lv et al., 2021), Freebase (Lv et al., 2021) and Yelp (Yang et al., 2020) to evaluate the performance of HERO. We adopt three widely used HGNN backbones in our experiments: HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), and HGT (Hu et al., 2020). See Appendix  B.1 for the details of the datasets and  B.4 for experiment setup.
Baselines
To evaluate the effectiveness of our proposed method, we select strong baselines from the Continual Graph Learning Benchmark (CGLB) (Zhang et al., 2022a), including Elastic Weight Consolidation (EWC) (Kirkpatrick et al., 2017), Memory Aware Synapses (MAS) (Aljundi et al., 2018), Experience Replay GNN (ER-GNN) (Zhou and Cao, 2021), and Topology-aware Weight Preserving (TWP) (Liu et al., 2021). EWC and MAS were previously not designed for graphs, so we train a HGNN on new tasks but ignore the graph structure when applying continual learning strategies. Additionally, we include two experience-replay-based baselines: MetaCLGraph (Unal et al., 2023) and FTF-ER (Pang et al., 2024). Furthermore, we use Finetune method (without any continual learning techniques) as a lower bound, and Joint Train (which allows access to previous data during training) as an approximate upper bound (Caruana, 1997).
Metrics
In our experiments, Average Performance (AP) and Average Forgetting (AF) (Lopez-Paz and Ranzato, 2017), are used to measure the performance on test sets. AP and AF are defined as $\text{AP}=\frac{1}{T}\sum_{t=1}^{T}a_{T,t},\,\,\,\,\text{AF}=\frac{1}{T-1}\sum_{t=1}^{T-1}\left(a_{T,t}-a_{t,t}\right),$ where $T$ is the total number of tasks and $a_{i,j}$} is the accuracy of the model on the test set of task $j$ after it is trained on task $i$.
Figure 3. Performance comparison on the Yelp dataset under two settings using HAN as the backbone model.
Performance Evaluation
The overall performance across three heterogeneous graph benchmarks is summarized in Table 1. Additional insights from task-wise accuracy retention on Yelp are shown in Figure 3 and Figure 4. We highlight the following key observations: (1) All HGNNs exhibit catastrophic forgetting on previous tasks. For example, on the DBLP dataset, the average forgetting (AF) for HAN and MAGNN is over 26% and 34%, respectively; on the IMDB dataset, the AF for all HGNNs is over 26%, 17%, and 21%, respectively. (2) Across all three HGNN backbones, HERO substantially outperforms state-of-the-art continual graph learning methods in terms of average performance (AP) while maintaining the lowest forgetting (AF). For example, on DBLP with HAN backbone, HERO improves AP by +2.7% over the strongest baseline (MAS) while reducing AF to 4.3%, close to the upper bound of joint training. Similar consistent performance gains are observed on IMDB and Freebase. These results confirm the effectiveness of HERO as a holistic framework for balancing knowledge adaptation and retention in HCGL. Furthermore, HERO demonstrates relatively low variance across the 5 runs, indicating its ability to consistently perform well in various scenarios. (3) It is noteworthy that, in some cases, memory-replay based methods (e.g., FTF-ER, MetaCLGraph, and ER-GNN) perform better than HERO in terms of AF, despite their lower AP scores compared to HERO. This is because these methods rely on retraining nodes from previously learned tasks. Once enough nodes are sampled, these models can largely mitigate catastrophic forgetting. However, this also sacrifices performance when learning new tasks. Although these methods attempt to address the new-old task trade-off through different strategies (e.g., ER-GNN reweights the loss between new and old tasks), they still cannot fully avoid the performance drop. We further discuss this in Appendix C.4. (4) We also observe a notable phenomenon: in the 3-way setting, memory-replay methods generally beat regularization-based methods, while in the 2-way setting the opposite holds. Further analysis of Figure 4 shows that regularization methods are less effective than memory-replay at overcoming forgetting of earlier tasks, but they achieve better performance on recent tasks, which is the “performance sacrifice” effect mentioned earlier. Memory-replay tends to accumulate redundant or conflicting information over long task sequences, which can harm adaptation to recent tasks. HERO mitigates this by filtering redundant information and applying knowledge distillation, preserving historical knowledge while maintaining stronger performance on recent tasks.
Figure 4. Visualization: Accuracy matrices on the Yelp dataset under 2-way setting.
Table 2. Ablation results of HERO using HAN as the backbone. We remove Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (G-MM) individually. And “w/o DiSCo” denotes replacing the proposed Diversity Sampling with semantic Consistency with Coverage Maximization applied to all node types.
Methods DBLP IMDB Freebase Yelp
AP / % $\uparrow$ AF / % $\downarrow$ AP / % $\uparrow$ AF / % $\downarrow$ AP / % $\uparrow$ AF / % $\downarrow$ AP / % $\uparrow$ AF / % $\downarrow$
HERO 93.6 $\pm$ 0.8 4.3 $\pm$ 1.6 78.4 $\pm$ 1.7 7.3 $\pm$ 3.0 60.9 $\pm$ 1.5 11.1 $\pm$ 1.6 55.9 $\pm$ 0.6 7.5 $\pm$ 1.2
w/o ER 89.7 $\pm$ 0.6 12.0 $\pm$ 0.7 75.1 $\pm$ 0.9 12.8 $\pm$ 1.7 58.6 $\pm$ 1.2 16.3 $\pm$ 2.2 53.9 $\pm$ 2.4 6.8 $\pm$ 2.4
w/o HKD 89.6 $\pm$ 1.4 12.1 $\pm$ 2.6 72.2 $\pm$ 1.5 19.3 $\pm$ 2.3 59.4 $\pm$ 1.1 14.7 $\pm$ 1.8 44.9 $\pm$ 14.3 19.3 $\pm$ 17.0
w/o G-MM 92.6 $\pm$ 1.7 5.2 $\pm$ 0.9 75.0 $\pm$ 2.7 13.3 $\pm$ 4.5 59.7 $\pm$ 0.9 12.0 $\pm$ 2.2 55.6 $\pm$ 1.9 8.0 $\pm$ 2.0
w/o DiSCo 92.4 $\pm$ 1.4 6.0 $\pm$ 3.1 75.4 $\pm$ 2.2 12.1 $\pm$ 5.0 56.2 $\pm$ 0.4 18.0 $\pm$ 0.5 45.7 $\pm$ 14.0 23.7 $\pm$ 15.6
Ablation Study
To evaluate the effectiveness of individual modules in HERO, we conducted comprehensive ablation experiments using HAN as the backbone network across four benchmark datasets. The results are presented in Table 2.
We first examined the impact of each component by sequentially removing Experience Replay (ER), Heterogeneity-aware Knowledge Distillation (HKD), and the Gradient-based Meta-learning Module (G-MM) while keeping other components intact. The results demonstrate that removing any module leads to performance degradation in both AP and AF, indicating each component’s essential role in our framework. Notably, the removal of G-MM shows relatively smaller performance impact (0.3-3.4% decrease in AP compared to 1.5-11.0% for ER/HKD), which aligns with its primary function of facilitating rapid adaptation to new tasks rather than long-term knowledge retention, the latter being mainly handled by the synergistic effect of ER and HKD. Particularly, the absence of HKD causes the most significant performance deterioration (3.6-12.0% increase in FR), highlighting its crucial role as the core mechanism against catastrophic forgetting.
Furthermore, we replaced our proposed Diversity Sampling with semantic Consistency (DiSCo) with the basic Coverage Maximization (CM) strategy. As shown in the last row of Table 2, DiSCo demonstrates substantial performance advantages, especially on edge-dense datasets like Yelp. These results confirm that DiSCo significantly outperforms conventional sampling methods in preserving the original topological structure of heterogeneous graphs. We further investigate the impact of each component in the HKD module in Appendix C.2.
Hyperparameters Analysis
We analyze the impact of key hyperparameters on model performance, including the loss weight ($\lambda_{kd}$) in knowledge distillation (KD), the sampling budget for experience replay, and the shot number (i.e., the number of sampled nodes used in meta-learning), as shown in Figure 5. We further discuss in Appendix C.3 the impact of the trade-off weight used to balance the losses of two submodules in the knowledge distillation component. The experimental findings are as follows: Distillation loss Weight ($\lambda_{kd}$): On IMDB, higher weights better support knowledge transfer, while lower values ($\lambda_{kd}<0.2$) lead to increased forgetting. On Freebase, smaller weights ($0.1\sim 0.3$) help mitigate forgetting, but larger weights ($\lambda_{kd}>0.3$) hinder adaptability. For Yelp, the 3-way setting introduces large task shifts, requiring higher weights for stability and effective forgetting mitigation. Sample Budget: The sample budget influences the stability-plasticity trade-off. Too few nodes lead to insufficient information, degrading performance, while excessive sampling limits the model’s adaptability to new tasks. As shown in Figure 5, on datasets like Yelp, a large sampling budget (e.g., ¿20) significantly impairs learning on new tasks due to the accumulation of historical information. Shot Number: The number of sampled nodes impacts the stability-plasticity trade-off. Meta-learning enables rapid adaptation to new tasks. As shown in the results, model performance steadily improves up to 10-shot, but declines at 20-shot, likely due to overfitting to the current task’s data patterns. The optimal shot number is 10, balancing historical information retention and computational efficiency.
Figure 5. Sensitivity analysis of HERO with respect to key hyperparameters. The shaded areas represent variances. From left to right: (1) distillation loss weight, (2) experience replay sampling budget, and (3) shot number used in meta-learning. Due to space constraints, we only report the Average Performance (AP) here.
Scalability Analysis
Based on the HAN backbone, we evaluate recent continual graph learning approaches and proposed HERO framework on four web graph datasets using a single RTX 3080Ti GPU. Table 3 reports the average training time per task. The results show that the regularization-based method TWP is clearly more efficient than replay-based methods in training, highlighting an open problem that we aim to address in future work. Nevertheless, our method significantly outperforms other replay-based approaches, demonstrating substantial improvements in computational efficiency over traditional replay-based methods. Although TWP is slightly more efficient than HERO, it consistently underperforms in terms of accuracy and forgetting across all benchmarks. Considering HERO’s strong empirical performance, we argue that HERO provides a practical balance between efficiency and effectiveness, which can also be effectively trained on large-scale and long-sequence HCGL tasks. Due to length limitation, we provide results of inference time in Appendix C.1.
Table 3. Training time (s) comparison on four datasets. Lower is better. Numbers in parentheses denote the ratio relative to HERO.
Method DBLP IMDB Freebase Yelp
TWP 0.1566 (x0.42) 0.0320 (x0.78) 0.0676 (x0.59) 3.0195 (x0.50)
FTF-ER 1.6193 (x4.39) 0.2382 (x5.78) 0.3514 (x3.06) 8.4238 (x1.39)
ER-GNN 0.4105 (x1.11) 0.0533 (x1.29) 0.0780 (x0.68) 6.6517 (x1.10)
MetaCLGraph 1.7702 (x4.80) 0.3101 (x7.53) 0.6144 (x5.34) 9.7515 (x1.61)
HERO (Ours) 0.3689 (x1.00) 0.0412 (x1.00) 0.1150 (x1.00) 6.0433 (x1.00)
6. Conclusion
In this work, we systematically study the problem of heterogeneous continual graph learning (HCGL), a critical yet underexplored challenge in dynamic web environments such as recommendation, knowledge graphs, and social networks. We present HERO, a heterogeneity-aware continual learning graph framework designed for evolving web graphs. HERO introduces two key innovations: (1) DiSCo (Diversity Sampling with semantic Consistency), a sampling strategy that maximizes node diversity and expands subgraphs along metapaths to preserve important semantic and structural information. This design enables HERO to achieve promising performance with a significantly smaller buffer size compared to existing replay-based approaches, highlighting its memory efficiency. (2) Heterogeneity-aware Distillation, which aligns semantics and representation across tasks, ensuring consistent knowledge transfer. By integrating these components within a holistic mechanism, HERO effectively balances stability and plasticity, achieving superior performance and scalability across multiple web graph benchmarks.
7. Acknowledgment
This project is supported by the National Science Foundation under CAREER Award IIS-2338878 and a generous research gift from Morgan Stanley.
Appendix A Extension to General HGNNs
Previously, we have discussed how to align semantic-level information via knowledge distillation in meta-path-guided attention networks. In this section, we further demonstrate that the proposed method can be easily extended to HGNNs without attention mechanisms (such as RGCN), enabling the preservation of semantic information from previous tasks.
For each pair of target-type nodes $v_{i},v_{j}\in V_{\tau_{t}}$, we define the attention score $e_{ij}$ based on their hidden representations from the penultimate layer of the network $h_{i}^{(L-1)}$ and $h_{j}^{(L-1)}$, using the weight matrix $W^{(L)}$ from the last layer:
(14) $e_{ij}=\left(h_{i}^{(L-1)}W^{(L)}\right)^{\top}\tanh\left(h_{j}^{(L-1)}W^{(L)}\right),$
We then apply softmax normalization to the attention scores for each node $v_{i}$ to obtain an attention vector:
(15) $\boldsymbol{\alpha}_{i}=\text{softmax}(e_{i1},e_{i2},\dots,e_{i|V_{\tau_{t}}|})$
We compute attention vectors for both the teacher model $\boldsymbol{\alpha}_{i}^{(T)}$ and the student model $\boldsymbol{\alpha}_{i}^{(S)}$, and define the semantic alignment loss as the $L_{2}$ distance between the two:
(16) $\mathcal{L}_{\text{sem}}=\sum_{i\in V_{\tau_{t}}}\left\|\boldsymbol{\alpha}_{i}^{(T)}-\boldsymbol{\alpha}_{i}^{(S)}\right\|_{2}$
In addition to the node similarity-based attention defined above, for edge-type-aware attention models such as HGT, attention is computed for each relation type $r\in\mathcal{R}$. Let $\boldsymbol{\beta}_{r}^{(T)}$ and $\boldsymbol{\beta}_{r}^{(S)}$ denote the attention vectors for relation $r$ from the teacher and student models, respectively. The semantic alignment loss for such edge-type-based attention mechanisms is formulated as:
(17) $\mathcal{L}_{\text{sem}}=\sum_{r\in\mathcal{R}}\left\|\boldsymbol{\beta}_{r}^{(T)}-\boldsymbol{\beta}_{r}^{(S)}\right\|_{2}$
This formulation ensures that the student model preserves relation-aware structural semantics by aligning its attention distribution with that of the teacher model on each edge type.
Appendix B Supplemental experiment setups
B.1. Details of dataset
DBLP is a bibliographic dataset in the field of computer science. A widely adopted subset is used, containing four research areas, where nodes represent authors, papers, terms, and venues.
IMDB is a dataset derived from a online movie information platform. A subset including classes such as Action, Comedy, Drama, Romance, and Thriller is selected for use.
Freebase is a large-scale knowledge graph originally constructed for web search and semantic applications. A subgraph comprising eight genres of entities and approximately one million edges is sampled following procedures outlined in prior studies.
Yelp is a review network constructed from businesses, users, locations, and reviews. Although node features are not available, many business nodes are annotated with one or more labels from sixteen predefined categories. As a prototypical web application graph, Yelp embodies real-world challenges in recommendation and user–business interaction modeling.
Table 4. Details of datasets and continual learning tasks setting
Node Classification #Nodes #Node Types #Edges #Edge Types #Classes
DBLP 26,128 4 239,566 6 4
IMDB 21,420 4 86,642 6 5
Freebase 180,098 8 1,057,688 36 7
Yelp 82,465 4 32,548,358 7 16
Figure 6. Illustration of the domain incremental setting.
B.2. Baseline Introduction
We introduce the baseline models that we choose to compare in the following section:
Elastic weight consolidation (EWC) (Kirkpatrick et al., 2017) is a technique that regularizes the loss function, such that the model is encouraged to only modify the weights that are less important for the previous tasks. This is achieved by penalizing changes to weights that have large importance for the previously learned tasks, thereby mitigating catastrophic forgetting when learning new tasks.
Memory aware synapses (MAS) (Aljundi et al., 2018) is a regularization-based method that measures the importance of parameters based on the sensitivity of predictions to these parameters. When learning a new task, the model adjusts the weights according to the network’s activations to update the important parameters relevant to the new task data.
Topology-aware weight preserving(TWP) (Liu et al., 2021) is a method introduced for graph continual learning that incorporates weight preservation mechanisms based on graph topology. By considering the local structure of the graph, TWP aims to overcome catastrophic forgetting. After computing the loss, the importance scores of the model weights are calculated according to the topology of the given graph, and the loss is regularized accordingly. This method leverages the properties of the graph.
ER-GNN (Zhou and Cao, 2021) stores samples from previous tasks as experiences and replays them when learning new tasks. After learning task $T_{i}$, sample nodes are saved to the buffer using a selection function. When learning the next task $T_{i+1}$, separate graphs are constructed for each learned task $\{T_{t}\}_{t=1,\dots,i}$. Then, the GNN is trained using these graphs. The overall loss is calculated by regularizing the current task’s loss with the loss from the separately constructed graphs through experience replay.
MetaCLGraph (Unal et al., 2023) combines experience replay and meta-learning. In this method, the initial parameters of the model are calculated using the current task data. When learning a new task, stored samples from previous tasks are merged with the current task data to form a new graph, which is then used to update the model parameters.
FTF-ER (Pang et al., 2024) combines feature and global topological information by normalizing and weightedly integrating two types of node importance scores to evaluate node importance. Specifically, it introduces the Hodge Potential Score (HPS) module to capture global topological information. When learning a new task, a subgraph is induced using all the experience nodes stored in the buffer, the overall loss is calculated as the sum of the loss for the current task and the loss of the subgraph.
B.3. Experimental Setting
Domain-Incremental Learning (Domain-IL) refers to a continual learning scenario where the task objective remains the same across tasks, but the data distribution (domain) shifts over time. This implies that the semantic meaning of the model’s output stays fixed. For example, in knowledge graph completion, each task may involve different sets of entities and relations, but the prediction goal—completing triplets—remains unchanged. Similarly, temporal data splits can form Domain-IL settings, where data from different time periods vary in distribution, yet the learning objective stays consistent.
B.4. Implementation Detail
To evaluate the effectiveness and generalizability of our method, we conduct experiments on three HGNNs backbones: HAN (Wang et al., 2019), MAGNN (Fu et al., 2020), and HGT (Hu et al., 2020). Adam optimizer is used and the initial learning rate is set to 0.005 for all datasets. For HAN, each task is trained for 200 epochs with an early stopping patience of 100. For MAGNN, each task is trained for 100 epochs, with the early stopping patience set to 5 on DBLP and 10 on IMDB. For HGT, all tasks are trained for 300 epochs with a patience of 30. For datasets trained with mini-batches, the batch size is set to 8. We adopt coverage maximization as the selection function for all experience-replay-based continual learning methods. The buffer size is uniformly set to 50 for the target node type in experience replay baselines (e.g., ER-GNN, MetaCLGraph, and FTF-ER), and 20 under mini-batch settings. For non-target node types, the buffer size is set to 200. The hyperparameters for regularization-based baselines (e.g., EWC, MAS, and TWP) are set following the TWP official repository and the benchmark study (Zhang et al., 2022a).
Our meta-learning module is configured under a 10-shot setting (2-way or 3-way depending on the dataset). For our method’s hyperparameters, the knowledge distillation loss weight $\lambda_{kd}$ is selected between 0.1 and 1.0 for different datasets, and the logit-level distillation weight $\lambda_{\text{logit}}$ is selected between 0.6 and 1.5. We generally set the experience replay loss weight $\lambda_{er}$ and the semantic-level distillation weight $\lambda_{\text{sem}}$ to 1.0 and 10.0, respectively. The temperature in knowledge distillation is globally set to 1.0. All experiments are repeated five times with random seeds on Nvidia RTX A4000 and RTX 4090D GPUs, and we report the mean and standard deviation across all methods and datasets.
Appendix C Additional Results and Analysis
C.1. Inference Time Comparison
We provide detailed inference time results in Table 5. We observe that HERO achieves higher efficiency than most baseline methods during the inference phase, further supporting its practicality in real-world deployment.
Table 5. Inference time (s) comparison on four datasets. HERO show the absolute inference time (s), while other methods are reported as relative difference (%). Positive difference means slower than HERO.
Method DBLP IMDB Freebase Yelp
TWP +1.99% +3.81% +2.22% -1.54%
FTF-ER +1.85% +2.98% +2.48% +0.01%
ER-GNN +1.02% +2.78% -1.13% -2.13%
MetaCLGraph +0.13% +3.54% +4.92% +0.80%
HERO (Ours) 3.4124 0.9815 2.5918 9.5556
C.2. Additional Ablation Study
We further analyze the contribution of the two submodules in the HKD module—logit-level distillation and semantic-level distillation—to overall model performance, and assess the effectiveness of our proposed design.
Figure 7 illustrates the impact of different distillation strategies—node-level only (“with node”), semantic-level only (“with sem”), and the combination of both (“with both”)—on the model’s Average Performance (AP) and Average Forgetting (AF) across four datasets (DBLP, IMDB, Freebase, Yelp). The results show that the combined strategy (“with both”) consistently achieves the highest AP and lowest AF across all datasets, demonstrating that logit-level and semantic-level distillation complement each other. Semantic-level distillation alone (“with sem”) exhibits weaker forgetting suppression on Freebase and Yelp, suggesting it may struggle to independently capture task-level semantic changes. On complex heterogeneous graphs such as Yelp, the synergy of both distillation levels yields especially significant performance improvements, indicating that addressing both structural and semantic forgetting is crucial.
Figure 7. Average performance (AP) and average forgetting (AF) of the HERO method under three settings: with logit-level distillation, with semantic-level distillation, and with both.
C.3. Additional Hyperparameters Analysis
We further analyze the impact of the weight $\lambda_{logit}$ used to balance the logit-level distillation loss in the Heterogeneity-aware Knowledge Distillation (HKD) module. Since $\lambda_{sem}$ is typically set to 10 in our experiments, we do not discuss it here.
Figure 8. Sensitivity analysis of the logit-level distillation weight $\lambda_{logit}$ in the HKD module. Left: Average Performance (AP); Right: Average Forgetting (AF).
Figure 8 shows the changes in AP (left) and AF (right) across datasets under varying logit-level distillation loss weights (x-axis). For most datasets, AP remains relatively stable, indicating robustness to this hyperparameter. IMDB and Yelp exhibit more fluctuation, likely due to significant distribution shifts across tasks, making them more sensitive to weight changes. The AF curves reveal that moderate weights (0.8 to 1.2) generally result in lower forgetting, while too high or too low weights cause an imbalance between new and old task focus. Yelp shows the smallest AF variation, suggesting that the HKD mechanism performs steadily in complex heterogeneous settings, aiding the model’s generalization ability.
Table 6. Forgetting-aware Gap (FaG) comparison on four datasets with HAN as the backbone model (averaged over 5 runs). Lower is better.
Method DBLP IMDB Freebase Yelp
Joint Train 0.69 2.47 11.59 11.58
EWC 2.15 2.78 10.17 1.62
MAS 0.40 0.97 9.53 2.09
TWP 1.03 2.27 10.44 5.58
ER-GNN 0.22 2.62 8.44 8.46
MetaCLGraph 1.36 2.87 13.68 13.61
FTF-ER 0.83 3.43 10.59 7.19
HERO (Ours) 0.63 1.31 9.90 6.88
C.4. Forgetting-aware Gap
AP (Average Performance) measures the mean test accuracy across all previously learned tasks, while AF (Average Forgetting) quantifies the performance drop on a specific task after learning subsequent ones. However, we observe that in order to retain knowledge from earlier tasks, the model often compromises its performance on the final task. To assess this trade-off, we introduce a new metric called Forgetting-aware Gap (FaG), defined as the difference between the test accuracy obtained by training the model solely on the final task and the test accuracy on the same task after completing all tasks under the continual learning framework:
(18) $\text{FaG}=a_{T}^{FT}-a_{T}^{CL}$
where $a_{T}$ denotes the test accuracy on the final task $\mathcal{T}_{T}$, with “FT” referring to the finetuning baseline (i.e., training only on task $\mathcal{T}_{T}$), and “CL” referring to the accuracy obtained after training with a continual learning method. This gap reflects the performance degradation caused by the inherent plasticity-stability trade-off in continual learning.
Results Analysis The experimental results indicate that the Joint Train method shows consistently large Final-task Performance Gaps (FaG) across all benchmark datasets, suggesting that directly replaying all data severely impairs adaptability to new tasks. MetaCLGraph and FTF-ER also suffer high FaG values on Freebase and Yelp (13.68 and 10.59), reflecting performance degradation under distribution shifts—a manifestation of the stability-plasticity dilemma. ER-GNN achieves the lowest FaG (0.22) on DBLP, showing good adaptability, but its FaG sharply rises to 8.46 on Yelp, indicating that simple replay mechanisms struggle with complex heterogeneous structures. In contrast, our HERO method consistently maintains lower FaG values. Notably, it achieves 1.31 on IMDB and 6.88 on Yelp, outperforming experience-replay and regularization-based methods like TWP (2.27 on IMDB). These results validate that HERO, through its meta-learning and heterogeneity-aware distillation design, effectively mitigates final-task degradation while preserving task adaptability. Overall, FaG serves as a useful metric to quantify the trade-off between knowledge retention and adaptation, further supporting the effectiveness of HERO in heterogeneous continual learning.